OpenStudy (anonymous):

Find the roots of 1/x-3-1/x+5=1/6 where x is not equal to 3,-5

Parth (parthkohli):

I think you can first get a common denominator.

Parth (parthkohli):

\[\dfrac{1}{x-3} - \dfrac{1}{x + 5} = \dfrac{x +5}{(x -3)(x +5)} + \dfrac{x - 3}{(x -3)(x + 5)}\]Proceed.